OpenStudy (anonymous):

A rectangular storage container with an open top is to have a volume of 10 m^3. The length of this base is twice the width. Material for the base costs $5 per square meter. Material for the sides costs $3 per square meter. Find the cost of materials for the cheapest such container.

ganeshie8 (ganeshie8):

what have you tried so far

OpenStudy (anonymous):

I don't even know how to approach it

ganeshie8 (ganeshie8):

this is a word problem, interpret the problem and always draw a rough diagram even if it is not necessary

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

next look at the given info : \(lwh = 10\) \(l = 2w\)

ganeshie8 (ganeshie8):

set up an equation for the cost and minimize it

ganeshie8 (ganeshie8):

see if u can write an expression for cost of making the total container

ganeshie8 (ganeshie8):

`Material for the base costs $5 per square meter. ` Area of base = \(lw \) so can we say the cost for making base is \(5lw\) ?

OpenStudy (anonymous):

I am so confused right now I'm sorry

ganeshie8 (ganeshie8):

its okay, word problems confuse me too go through the question again and see if the problem makes sense first

OpenStudy (anonymous):

I am just confused with the money aspect of it

ganeshie8 (ganeshie8):

ikr, thats the most tricky part lets see

ganeshie8 (ganeshie8):

did you get why the cost of making base equals \(5*lw\) ?

OpenStudy (anonymous):

Yes I got that part

ganeshie8 (ganeshie8):

then we are good :) lets keep going

ganeshie8 (ganeshie8):

lets find the cost of making sides

ganeshie8 (ganeshie8):

`Material for the sides costs $3 per square meter. ` lateral area of box = (perimeter of base)*height = \(2(l+w)*h\) multiply it by 3 to get the total cost : \(3 *2(l+w)*h\)

ganeshie8 (ganeshie8):

so far we have this : cost of making base = \(5lw\) cost of making sides = \(3*2(l+w)*h\)

ganeshie8 (ganeshie8):

So our cost expression for total cost of making the box would be : \[C = 5lw + 3*2(l+w)*h\]

ganeshie8 (ganeshie8):

see if that makes sense so far

OpenStudy (anonymous):

It does.

ganeshie8 (ganeshie8):

good, we are not done yet we need to find the minimum value of that Cost function

ganeshie8 (ganeshie8):

you must be knowing how to minimize a function having single variable

ganeshie8 (ganeshie8):

but in our cost function we have 3 variables, what to do

ganeshie8 (ganeshie8):

somehow we need to eliminate 2 variables

OpenStudy (anonymous):

l can equal 2w right?

ganeshie8 (ganeshie8):

Yes! we can use the given info replace \(l\) by \(2w\)

ganeshie8 (ganeshie8):

\[C = 5lw + 3*2(l+w)*h\] becomes \[C = 5(2w)w + 3*2(2w+w)*h\]

ganeshie8 (ganeshie8):

still we have 2 variables we need to eliminate h somehow..

OpenStudy (anonymous):

I have no idea with that one.. Sorry

ganeshie8 (ganeshie8):

use the other given info

ganeshie8 (ganeshie8):

\(lwh = 10\) isolate h from here and plug it in ur cost function

OpenStudy (anonymous):

okay so \[c=5(2w)w+3*2(2w+w)*10/2w^2\]

ganeshie8 (ganeshie8):

Excellent! simplify a bit before starting the calculus part

ganeshie8 (ganeshie8):

simiplifying should give you \[C=10w^2+\frac{90}{w}\] check if it is correct ^

OpenStudy (anonymous):

That is what I got too

ganeshie8 (ganeshie8):

great! go ahead and minimize it

ganeshie8 (ganeshie8):

familiar with the process ? taking the first derivative and setting it equal to 0..

OpenStudy (anonymous):

Yes I got that part. It's the easiest

ganeshie8 (ganeshie8):

I agree, calculus is the easiest part :) algebra sucks though

ganeshie8 (ganeshie8):

wolfram gives a weird number for cost http://www.wolframalpha.com/input/?i=minimize+10w%5E2+%2B+90%2Fw

ganeshie8 (ganeshie8):

\[\large \text{minimum cost} = 45 \sqrt[3]{6}\]

OpenStudy (anonymous):

I got \[\sqrt[3]{9/20}\]

ganeshie8 (ganeshie8):

you got that for \(w\) right ?

OpenStudy (anonymous):

Yes

ganeshie8 (ganeshie8):

you should get \(\large w = \sqrt[3]{\frac{9}{2}}\)

ganeshie8 (ganeshie8):

check ur work again

OpenStudy (anonymous):

I found out what I did wrong